A proposed project has fixed costs of $82,000 per year. The operating cash flow at 6,200 units is $88,400. Ignoring the effect of taxes, what is the degree of operating leverage? (Do not round intermediate calculations. Round your answer to 4 decimal places, e.g., 32.1616.) Degree of operating leverage If units sold rise from 6,200 to 6,700, what will be the new operating cash flow? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.) Operating cash flow $ What is the new degree of operating leverage? (Do not round intermediate calculations. Round your answer to 4 decimal places, e.g., 32.1616.) New DOL
Answer #1

a. The Degree of operating leverage, DOL is given by:-

DOL = 1 + Fixed costs/Operating cash flows

DOL = 1 + $82,000/$88,400

DOL = 1 + 0.9276

DOL= 1.9276

b. Percentage change in quantity sold = (6700-6200)/6200

Percentage change in quantity sold = 8.06%

Percentage change in operating cash flows = DOL or 1.9276 x (8.06%)

Percentage change in operating cash flows = 15.53%

New operating cash flows = [(1 + % of changes in operating cash flows) x old operating cash flows]

New operating cash flows = (1 + 0.1553) x 88400

New operating cash flows = $102128.52

Degree of operating leverage = 1 + fixed cost / new operating cash flows

Degree of operating leverage = 1 + 82000/102128.52

Degree of operating leverage = 1.8029

Note: Question not formatted properly.